OpenStudy (anonymous):

An ice cream cone is 5" deep and 2" in top diameter. One scoop of vanilla ice cream, 2" in diameter, is placed on the cone. If the ice cream melts into the cone, will it overflow?

OpenStudy (anonymous):

Vcone= (pi*1²*5)/3 = 5.24 Vscoop= 4*(pi*1³)/3 = 4.19 So... Vcone>Vscoop ---> It won't overflow!